radomes.openskiesproject.org
Mh5cerqKeafitdroul
Mh5cerqKeafitdroul Mh5cerqKeafitdroul
  • 21-06-2016
  • Mathematics
Answered

The probability that Aaron will win the drawing is 0.083. What is the probability that he will not win?



0.17



0.083



0.83



0.917

Verified Answer :

Anonymous Anonymous
  • 21-06-2016

Let the probability that he will loose be x

x + 0.083 = 1

x = 1-0.083

x = 0.917

so 0.917 is the answer.

Answer Link
Anonymous Anonymous
  • 21-03-2019

Answer: 0.917

Step-by-step explanation: 1 - 0.083 = 0.917

Answer Link

Related Questions

What was the main purpose of war in Aztec society to appease the Aztec gods to acquire captives for sacrifice to gain territory to try out new weapons
Why are both accuracy and reproducibility important in making a measurement
Which display shows a bivariate data set
bravery plays an Important role in the story which details suggest oona is brave in this part of the story
Maya thinks that the theater sold 1256 regular price tickets Jonah says that 1248 regular price tickets were sold do you agree with either person explain your r
Gail would like to save for her childs college expenses to begin in 18 years Which of the following methods would be the easiest way for Gail to save for over 1
What is it when blood flows through two consecutive capillary networks before returning to the heart a Serial circulation b Parallel circulation c Collateral ci
Help me plzzzzzxxxx
The Suarez family paid 1650 for 3 movie tickets How much would they have paid for 9 tickets
will upvote Which of the following best describes how Hammurabis Code and the Jewish Ten Commandments are similar A They established a public and consistent set